OpenStudy (anonymous):

Which polynomial is a perfect square trinomial? (4 points) A 25x2 - 20x - 16 B 9a2 - 30a - 25 C 25b2 - 15b + 9 D 16x2 - 24x + 9

OpenStudy (solomonzelman):

which options can you exclude right away ?

OpenStudy (anonymous):

erm... D?

OpenStudy (solomonzelman):

The last term, when factoring ANY expression \(\large\color{slate}{ (x-a)^2 }\) should [always] be negative or positive ?

OpenStudy (solomonzelman):

\(\large\color{slate}{ (x-a)^2~~\Rightarrow~~x^2-2ax+a^2 }\)

OpenStudy (anonymous):

positive

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

(and certainly \(\large\color{slate}{ (x+a)^2 }\) wouldn't have any negative -- which implies for sure no negative last term)

OpenStudy (solomonzelman):

So which options do you have left /

OpenStudy (anonymous):

so A and B are thrown out.

OpenStudy (anonymous):

c and d

OpenStudy (solomonzelman):

yes, very good!

OpenStudy (solomonzelman):

C. \(\large\color{slate}{ 25b^2 - 15b + 9 }\) D. \(\large\color{slate}{ 16x^2 - 24x + 9 }\)

OpenStudy (anonymous):

C?

OpenStudy (solomonzelman):

always keep in mind: \(\large\color{blue}{ (A-B) ^2=A^2-2AB+B^2}\) For C, \(\large\color{slate}{ 25b^2 - 15b + 9 }\) \(\large\color{slate}{ (5b)^2 - 15b + 9 }\) \(\large\color{slate}{ (5b)^2 - 15b + (3)^2 }\) so for this to be a perfect square, you must be having: (look at the blue again) \(\large\color{slate}{ (5b)^2 - 15b + (3)^2~~~\Rightarrow~~\color{blue}{ (5b-3)^2} }\) Now, check if this (last blue expression (5b-3)^2 obtains the middle term. For D, \(\large\color{slate}{ 16x^2 - 24x + 9 }\) \(\large\color{slate}{ (4x)^2 - 24x + 9 }\) \(\large\color{slate}{ (4x)^2 - 24x + (3)^2 }\) so for this to be a perfect square, you must be having: (look at the blue again) \(\large\color{slate}{ (4x)^2 - 24x + (3)^2~~~\Rightarrow~~\color{blue}{ (4x-3)^2} }\) Now, check if this (last blue expression (4x-3)^2 obtains the middle term.

OpenStudy (anonymous):

I still think C

OpenStudy (solomonzelman):

AGAIN, \(\large\color{blue}{ (A-B) ^2=A^2-2AB+B^2}\) So for your middle term to be 15, the following has to be true. \(\large\color{slate}{ (5)(3)(2)=15~? }\)

OpenStudy (solomonzelman):

see the fault here ?

OpenStudy (anonymous):

oh.. i do. So the correct answer is D! thank you!

OpenStudy (solomonzelman):

yes, you could have reduced to D using elimination

OpenStudy (solomonzelman):

I will elaborater

OpenStudy (solomonzelman):

\(\large\color{slate}{ (a-b)^2=a^2-2ab+b^2 }\) (we got that piece) So if \(\large\color{slate}{ a }\) and \(\large\color{slate}{ b }\) are integers (as we had in all cases) then, will the red part (i.e the middle term) \(\large\color{slate}{ (a-b)^2=a^2-\color{red}{2ab}+b^2 }\) be ever an odd number ?

OpenStudy (solomonzelman):

integers, btw are \(\large\color{slate}{ {\bf...}~-3,~-2,~-1,~0,~~1,~~2,~~3~{\bf...} }\) all those positive and negative "whole" numbers.

OpenStudy (solomonzelman):

trouble(s), concern(s), question(s) ?